Good day, Can you please advise on a GRAPH for EQUILIBRIUM quantity calculation of Qs=18000+0.2P & QUANTITY DEMANDED Calcultuon Qd=2400-0.1P


1
Expert's answer
2021-03-31T17:06:29-0400

The market equilibrium occurs when "Q_d=Q_s":


"2400-0.1P_E=1800+0.2P_E,""0.3P_E=600,""P_E=\$2000."

Then, we can substitute "P_E" into the demand function and find the equilibrium quantity:


"Q_E=2400-0.1\cdot\$2000=2200."

Let's draw a graph for the market equilibrium: